A solenoid of length 2.2 m has a diameter of
0.9 m. The magnetic field at its centre is
0.4 tesla. Estimate the energy stored in the
field of this coil
1
Expert's answer
2020-08-14T15:57:43-0400

Given

Length of coil is "l=2.2 m"

Magnetic field at the center of the solenoid is "B=0.4 T"

The energy stored in the solenoid is

"U=\frac{B^2Al}{2\mu_0}=\frac{(0.4)^2(\pi(0.45)^2)(2.2)}{2(4\pi\times10^{-7} )}"

"=8.91\times10^4 T"